40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 15 16:46:27 2021, Anchal, Summary, Optical Levers, Vent Prep Screen_Shot_2021-10-15_at_4.34.45_PM.pngScreen_Shot_2021-10-15_at_4.45.31_PM.pngScreen_Shot_2021-10-15_at_4.34.45_PM.pngScreen_Shot_2021-10-15_at_4.34.45_PM.png
    Reply  Fri Oct 15 17:17:51 2021, Koji, Summary, General, Vent Prep 7x
       Reply  Fri Oct 15 20:53:49 2021, Koji, Summary, General, Vent Prep 
Message ID: 16407     Entry time: Fri Oct 15 16:46:27 2021     Reply to this: 16408
Author: Anchal 
Type: Summary 
Category: Optical Levers 
Subject: Vent Prep 

I centered all the optical levers on ITMX, ITMY, ETMX, ETMY, and BS to a position where the single arm lock on both were best aligned. Unfortunately, we are seeing the TRX at 0.78 and TRY at 0.76 at the most aligned positions. It seems less power is getting out of PMC since last month. (Attachment 1).

Then, I tried to lock PRMI with carrier with no luck. But I was able to see flashing of up to 4000 counts in POP_DC. At this position, I centered the PRM optical lever too (Attachment 2).

Attachment 1: Screen_Shot_2021-10-15_at_4.34.45_PM.png  85 kB  Uploaded Fri Oct 15 17:48:25 2021  | Hide | Hide all
Screen_Shot_2021-10-15_at_4.34.45_PM.png
Attachment 2: Screen_Shot_2021-10-15_at_4.45.31_PM.png  18 kB  Uploaded Fri Oct 15 17:50:05 2021  | Hide | Hide all
Screen_Shot_2021-10-15_at_4.45.31_PM.png
Attachment 3: Screen_Shot_2021-10-15_at_4.34.45_PM.png  85 kB  Uploaded Fri Oct 15 17:50:45 2021  | Hide | Hide all
Screen_Shot_2021-10-15_at_4.34.45_PM.png
Attachment 4: Screen_Shot_2021-10-15_at_4.34.45_PM.png  85 kB  Uploaded Fri Oct 15 17:51:01 2021  | Hide | Hide all
Screen_Shot_2021-10-15_at_4.34.45_PM.png
ELOG V3.1.3-